OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 4.

OpenStudy (anonymous):

the vertex of the graph is -b/2a

OpenStudy (anonymous):

a = -3 b = 6 c = -4

OpenStudy (anonymous):

-6/-6

OpenStudy (anonymous):

vertex=1

OpenStudy (anonymous):

no that is the x now you got to plug it into the x and find y

OpenStudy (cwrw238):

to be more exact the x coordinate of th vertex = -b/2a

OpenStudy (anonymous):

so the vertex is (1,-1)

OpenStudy (anonymous):

Can u do another

OpenStudy (anonymous):

Identify the vertex for the graph of y = 2x2 + 8x − 3.

OpenStudy (anonymous):

@cwrw238 you can do this one

OpenStudy (anonymous):

ok let me try

OpenStudy (cwrw238):

best if you try yourself

OpenStudy (anonymous):

-4,4

OpenStudy (cwrw238):

-b/2a = -8 / 4 = -2 x is -2 and y = 2(-2)^2 -16 -3

OpenStudy (cwrw238):

so that (-2,-11)

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

another way to do these is to convert the equation into vertex form 2x2 + 8x − 3. = 2(x"2 + 4) - 3 = 2(x + 2)^2 _ 8 - 3 = 2*x + 2)^2 - 11 the x value of the vertex is when x+2 = 0 and the y value is -11 = (-2,-11)

OpenStudy (cwrw238):

* = 2(x + 2)^2 - 11